A bag has 7 red, 2 blue and 3 green balls. One ball is picked at random. P( not red) =
- 512
- 712
- 57
- 16
Answer
A) 512
P( not red) = 1 - P( red) = 1 - 712 = 512. SLO M-09-C-04 · Knowledge
Mathematics MCQ · More Probability MCQs · Practise
A) 512
P( not red) = 1 - P( red) = 1 - 712 = 512. SLO M-09-C-04 · Knowledge
Mathematics MCQ · More Probability MCQs · Practise